OpenStudy (anonymous):

A ball is thrown vertically upward from the top of a 100-foot tower, with an initial velocity of 20 ft/sec. Its position function is s(t) = –16t2 + 20t + 100. What is its velocity in ft/sec when t = 1 second?

OpenStudy (anonymous):

-12 –44 100 –32

OpenStudy (anonymous):

-16t^2

OpenStudy (anonymous):

@Compassionate @Aamirgsen @cwrw238 @iambatman @Callisto @uri @eliassaab @hba @PRAETORIAN.10

OpenStudy (cwrw238):

do you know how to find the formula for the velocity ? Can you find ds/dt?

OpenStudy (cwrw238):

this is calculus.

OpenStudy (anonymous):

to find velocity we do just find the derivative

OpenStudy (anonymous):

right

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

i get -32t + 20 i just dont know wht do do afterwards

OpenStudy (anonymous):

yes the derivative is correct. then set t=1

OpenStudy (anonymous):

because you need the velocity at 1 second (at t=1)

OpenStudy (anonymous):

so -32(1)+12=-12 is tht right then wht else do i do

OpenStudy (anonymous):

nothing you are done.

OpenStudy (anonymous):

you found the velocity when t=1

OpenStudy (anonymous):

how r u sure

OpenStudy (cwrw238):

fakee is correct -12

OpenStudy (anonymous):

okay...cool can you guys help me with another problem?

OpenStudy (anonymous):

you have a position function, they ask you to find the velocity at t=1. so you find the veloicty function, (which is s'(t), the derivative) and plug in t=1. basically, you were asked s'(1).

OpenStudy (anonymous):

What is the instantaneous slope of y=8/x at x=-3

OpenStudy (anonymous):

f(x)=8/x instantaneous rate at x=a, is f'(a)

OpenStudy (anonymous):

so 8

OpenStudy (anonymous):

f '(a), I said:)

OpenStudy (anonymous):

i meant -3

OpenStudy (anonymous):

find the derivative fisrt

OpenStudy (anonymous):

first*

OpenStudy (anonymous):

which is 1/x

OpenStudy (anonymous):

no

OpenStudy (cwrw238):

no - do you know what the derivative of 1/x is?

OpenStudy (anonymous):

1/x is same as x^-1.

OpenStudy (anonymous):

use the power rule.

OpenStudy (anonymous):

its 1

OpenStudy (anonymous):

\[\frac{d}{dx} x^n= nx^{x-1}\]

OpenStudy (anonymous):

8 is a k so 0???

OpenStudy (anonymous):

\[\frac{d}{dx} x^{-1}= ?\]

OpenStudy (anonymous):

8

OpenStudy (anonymous):

i dont know how to do thios

OpenStudy (anonymous):

this

OpenStudy (anonymous):

the power rule:\[\frac{d}{dx} x^n= nx^{x-1}\]your question:\[\frac{d}{dx} x^{\color{red}{-1}}= ({\color{red}{-1}})x^{{\color{red}{-1}~-1}}\]

OpenStudy (anonymous):

what do you get for the derivative? (I am using d/dx notation for the derivative)

OpenStudy (anonymous):

solving

OpenStudy (anonymous):

and you get? or still lost?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

when you did

OpenStudy (anonymous):